A quiz consists of 20 true or false questions. If the student guesses on each question, what is the Variance and the standard deviation of the number of
correct answers?

Answer by ewatrrr(24785)


Hi
20 t or false questions guessed.  p(correct answer) =  1/2 
mean =  20(1/2) = 10
variance = 10(1-p) = 10(1/2) = 5
sd = √5 =    

Wish You the Best in your Studies.